OpenStudy (anonymous):

Emily has 3 yellow, 8 green, 5 red, and 14 blue ankle bracelets in a bag. If Emily randomly selects an ankle bracelet out of the bag, what is the probability it will be green or blue?

OpenStudy (dumbcow):

total = 30 num green = 8 num blue = 14 P(GorB) = (8+14)/30 = 22/30 = 11/15

OpenStudy (anonymous):

These sort of probability problems are quite easy. All you have to do is find the total number of ankle bracelets. Which is 30 total ankle bracelets. Next you need to find the total number of colored bracelets you could possibly pull out of the bag. In this case you are working with green and blue. So what you need to do is add the number of blue and green bracelets together. This makes 14+8=22. Now you place 22 as your numerator and 30 as your denominator to make 22/30. Now I am not sure if you have learned to simplify or if the problem asks but just in case I will. The simplified answer is 11/15. Hope this helps!